# Right-handed neutrino dark matter consistent with the generation of baryon number asymmetry

### v1

@@ 2.2 Relic abundance of right-handed neutrino DM
Fig. 1   Left panel: Evolution of the number density of $N_{1}$ and $\eta_{R}$ in the comoving volume for three values of $M_{1}$ are plotted with red lines, 1.5 TeV (solid), 150 GeV (dashed) and 1.5 GeV (dotted). $h_{1}$ is fixed at $7.8\times 10^{-9}$, $4.5\times 10^{-8}$ and $7.0\times 10^{-6}$ for each $M_{1}$, respectively. $M_{\eta_{R}}=3$ TeV and $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$ are assumed. Horizontal black lines represent $Y_{N_{1}}$ values required for each value of $M_{1}$ to realize the present DM relic density. While an $Y_{\eta_{R}}$ line for $M_{1}=1.5$ TeV deviates from $Y_{\eta_{R}}^{\rm eq}$, $Y_{\eta_{R}}$ in other two cases coincides with $Y_{\eta_{R}}^{\rm eq}$ since the $\eta_{R}$ decay is already in the equilibrium at a temperature higher than $T_{c}$. Right panel: Contours of $Y_{\eta_{R}}(T_{\eta})$ for $M_{\eta_{R}}=3$ TeV in the $(\lambda_{3},\lambda_{4})$ plane. Each red line represents $Y_{\eta_{R}}(T_{\eta})=10^{-10.5},10^{-11},10^{-11.5},10^{-11.75}$, and $10^{-12}$ outward. A dashed line represents $\lambda_{3}+\lambda_{4}+\lambda_{5}=-\sqrt{\lambda_{1}\lambda_{2}}$ in the case of $\lambda_{2}=1.0$, and an upper region of it satisfies the potential stability. No constraint comes from direct detection experiments on $\lambda_{3,4}$ since DM is $N_{1}$.
@@
In order to show this observation concretely, we solve the Boltzmann equations for $Y_{N_{1}}$ and $Y_{\eta_{R}}$ using the parameters given in eq. (14) and $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$. Since $M_{1}$ and $h_{1}$ can be free from neutrino mass determination as addressed below eq. (14), we plot the solutions in the left panel for the cases $(M_{1},h_{1})=(1.5~{}{\rm TeV},7.8\times 10^{-9})$, $(150~{}{\rm GeV},4.5\times 10^{-8})$, and $(1.5~{}{\rm GeV},7.0\times 10^{-6})$ where $h_{1}$ is tuned to give the required DM abundance. In this panel, $Y_{N_{1}}$ required to explain the DM abundance is shown by horizontal black lines for each case. In the latter two cases, the required $Y_{N_{1}}^{\infty}$ is realized through the pure freezein due to the $\eta_{R}$ decay. $Y_{\eta_{R}}$ does not deviate from the equilibrium one since its decay to $N_{1}$ is already in the equilibrium at $T>T_{\eta}$ and then $Y_{\eta_{R}}=Y_{\eta_{R}}^{\rm eq}$ is kept. The freezeout of the $\eta_{R}$ coannihilation in a case of $M_{\eta_{R}}=3$ TeV is shown in the right panel, in which contours for $Y_{\eta_{R}}^{\rm eq}(T_{\eta})=10^{-10.5},10^{-11},10^{-11.5},10^{-11.75}$, and $10^{-12}$ are plotted in the $(\lambda_{3},\lambda_{4})$ plane. Since $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$ is assumed in the left panel, we find $Y_{\eta_{R}}^{\rm eq}(T_{\eta})=10^{-11.75}$ is relevant. It is smaller than the required $Y_{N_{1}}^{\infty}$ in these cases, and then the freezeout of the $\eta_{R}$ coannihilation is irrelevant to the estimation of $Y_{N_{1}}^{\infty}$. As an extreme case of these examples, we may consider $N_{1}$ with keV scale mass which can be consistent with the neutrino oscillation data as discussed in the previous part. We can confirm that the required relic DM abundance is obtained by tuning $h_{1}$ suitably, for example, $h_{1}=3.7\times 10^{-3}$ for $M_{1}=5$ keV.
@@
On the other hand, the freezeout of the $\eta_{R}$ coannihilation is relevant to the estimation of $Y_{N_{1}}$ in the case of $M_{1}=1.5$ TeV. Although $Y_{N_{1}}(T_{c})$ is smaller than the required value, the late time decay of the frozen $\eta_{R}$ contributes to $Y_{N_{1}}$ and make it the required value $Y_{N_{1}}^{\infty}$ as found from the figure. Since the stability of the potential $V$ in eq. (1) requires $\lambda_{3,4}$ should satisfy the condition (2), an allowed region of $(\lambda_{3},\lambda_{4})$ is a part over the dashed line in the right panel. Although $Y_{N_{1}}$ is determined by the values of $\lambda_{3}$ and $\lambda_{4}$ in the $M_{1}=1.5$ TeV case, we note that there is no constraint on them from direct detection experiments since $N_{1}$ is DM which cannot interact with nuclei. It is a crucial difference from the case where $\eta_{R}$ is DM. The coupling constants $\lambda_{3}$ and $\lambda_{4}$ could be also constrained through the correction to the Higgs mass caused by them. This constraint is reported to be satisfied for $M_{\eta_{R}}\leq 4.4$ TeV in [27].
@@
In the above study, we do not take into account the effect caused by the simultaneous explanation of the baryon number asymmetry in the Universe through leptogenesis based on the $N_{2}$ decay. Since it could produce $\eta_{R}$ at late time, the $N_{1}$ abundance could be affected by it if the $N_{2}$ decay occurs near the freezeout of $\eta_{R}$. We discuss this point in the next part.
@@ 3 Consistency with leptogenesis
In the present model, the baryon number asymmetry is considered to be produced through leptogenesis caused by the out-of-equilibrium decay of $N_{2}$. This requires that $N_{2}$ should be in the thermal bath and its decay to $\eta_{R}$ occurs in a $CP$ asymmetric way at a temperature higher than 100 GeV where sphaleron interaction is in the equilibrium. Values of $h_{2,3}$ given in eq. (14) constrained by the neutrino oscillation data are large enough to produce $N_{2,3}$ in the thermal bath as relativistic particles through the inverse decay and scatterings as long as reheating temperature is larger than $M_{3}$. $CP$ asymmetry in the decay $N_{2}\rightarrow\ell_{\alpha}\eta$ can be caused through an interference between a tree diagram and a one-loop diagram mediated by $N_{1}$ and $N_{3}$.
The $CP$ asymmetry caused by one loop vertex diagrams is estimated through the formula
@@
where $M_{2}\simeq M_{\eta_{R}}$ is taken into account. If we note that $h_{1}$ is supposed to be very small for the DM explanation as discussed in the previous part, the dominant contribution to this $\varepsilon$ is found to be caused by $N_{3}$ and it can be estimated by using eqs. (7) and (12) asdddIt should be noted that the existence of both a zero mass eigenvalue and Majorana phases of the right-handed neutrino masses are required to guarantee a non-zero $\varepsilon$ when neutrino Yukawa couplings $h_{\alpha k}$ are real [14].
@@
where parameters in eq. (14) and $\sin\theta_{ij}$ and $\alpha_{ij}$ written below it are used, and $\sin(\gamma_{2}-\gamma_{3})=1$ is assumed in the second line. We should note that a smaller $|\lambda_{5}|$ can make Yukawa couplings larger and then $\varepsilon$ larger but it makes the out-of-equilibrium decay of $N_{2}$ harder.
The $N_{2}$ decay is in the equilibrium at temperature $T\geq T_{D}$. $T_{D}$ is defined by $\Gamma_{N_{2}}=H(T_{D})$ where the decay width of $N_{2}$ can be expressed as
@@
If the washout processes of the generated lepton number asymmetry are frozen out at temperature $T_{F}$, the lepton number asymmetry caused by the out-of-equilibrium decay of $N_{2}$ could escape the washout in the case $T_{D}<T_{F}$ since the dominant part of lepton number asymmetry is expected to be generated at $T~{}{{}^{<}_{\sim}}~{}T_{D}$. Such situation could happen if the $N_{2}$ decay is forced to be delayed and be kept in the out-of-equilibrium by making $T_{D}$ to a lower temperature. On the other hand, the out-of-equilibrium decay of $N_{2}$ is favored to start at the temperature $T~{}{{}^{>}_{\sim}}~{}M_{2}$ for the efficient generation of the lepton number asymmetry since the $N_{2}$ abundance can keep a large value as the relativistic particle there.
Although $h_{2}$ fixed by the neutrino oscillation data cannot realize such a situation for a TeV scale $M_{2}$, it could be realized through an additional suppression factor for $\Gamma_{N_{2}}$. It is caused the phase space suppression due to the strict degeneracy between $M_{2}$ and $M_{\eta_{R}}$ [16]. In fact, if their degeneracy is expressed as $M_{\eta_{R}}=(1-\delta_{2})M_{2}$ with $\delta_{2}\ll 1$, $\Gamma_{N_{2}}$ is found to be rewritten with an additional suppression factor $\delta_{2}$ as
$\Gamma_{N_{2}}\simeq\frac{h_{2}^{2}\delta_{2}^{2}}{4\pi}M_{2}=2.9\times 10^{-12}\left(\frac{M_{2}}{3~{}{\rm TeV}}\right)\left(\frac{\delta_{2}}{10^{-4}}\right)^{2}~{}{\rm GeV},$ (24)
where we use the values for $\sin\theta_{ij}$ and $\alpha_{ij}$ given below eq. (14). Since it shows that $\Gamma_{N_{2}}<H(M_{2})$ is satisfied for $\delta_{2}<10^{-4}$ at $M_{2}=3$ TeV, $N_{2}$ is expected to be out-of-equilibrium at $T~{}{{}^{>}_{\sim}}~{}M_{2}$. Thus, if the condition $T_{D}<T_{F}$ is fulfilled, the generated lepton number asymmetry $Y_{L}$ defined as $Y_{L}\equiv Y_{\ell}-Y_{\bar{\ell}}$ could be estimated as $Y_{L}=\varepsilon Y_{N_{2}}^{\rm eq}$ where $Y_{N_{2}}^{\rm eq}$ stands for the equilibrium value of $N_{2}$ as the relativistic particle. In that case, lepton number asymmetry $Y_{L}=O(10^{-10})$ required for the explanation of baryon number asymmetry in the Universe is expected to be obtained if $\varepsilon$ takes a value of $O(10^{-7})$ in eq. (22). Unfortunately, since $\delta_{2}\ll 1$ also suppresses $\varepsilon$ heavily as found from eq.(21), this scenario cannot work.
@@
It is crucial to know what value of $\delta_{2}$ can make the condition $100~{}{\rm GeV}<T_{D}<T_{F}$ be satisfied for the parameters given in eq. (14). By calculating the relevant reaction rates of washout processes, we find that $T_{F}~{}{{}^{>}_{\sim}}~{}0.1M_{2}$ so that $\delta_{2}<3\times 10^{-5}$ guarantees $T_{D}<T_{F}$. This suggests that the $CP$ asymmetry $\varepsilon$ can take a value of $O(10^{-7})$ in the case $\delta_{3}=O(10^{-9})$, which makes successful leptogenesis possible as addressed below eq. (24). However, if the number density of $N_{2}$ does not decrease to the required $Y_{N_{1}}^{\infty}$, the $\eta_{R}$ produced through the $N_{2}$ decay contributes to the $N_{1}$ relic density. Thus, the estimation of the $N_{1}$ relic abundance in the previous section cannot naively applied to this scheme for leptogenesis. The effect of $\eta_{R}$ produced in the late time $N_{2}$ decay needs to be taken into account for the estimation of the $N_{1}$ relic abundance.
@@
$\frac{dY_{N_{2}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left\{\gamma^{S}(N_{2}\ell)+\gamma^{S}(N_{2}\eta)+\gamma^{D}(N_{2})\right\}\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right),$
$\frac{dY_{L}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\varepsilon\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)+\left\{\sum_{k=2,3}\gamma^{D}(N_{k})+2(\gamma^{S}(\ell\eta)+\gamma^{S}(\ell\ell))\right\}\frac{Y_{L}}{Y_{\ell}^{\rm eq}}\right],$
$\frac{dY_{\eta_{R}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\left\{2\gamma^{S}(\eta_{R}\eta_{R})+\gamma^{D}(\eta_{R})\right\}\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right)+\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)\right],$
$\frac{dY_{N_{1}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})}\gamma^{D}(\eta_{R})\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right),$ (27)
@@
where $\gamma^{S}({N_{2}\ell})$ and $\gamma^{S}(N_{2}\eta)$ are reaction densities of the scattering $N_{3}\ell\leftrightarrow N_{2}\ell$ and $N_{3}\eta\leftrightarrow N_{2}\eta$ for $N_{2}$ production, and $\gamma^{S}(\ell\eta)$ and $\gamma^{S}(\ell\ell)$ are the ones for the lepton number violating scattering $\ell\eta\leftrightarrow\bar{\ell}\eta^{\dagger}$ and $\ell\ell\leftrightarrow\eta\eta$. A new term is introduced to the equation for $Y_{\eta_{R}}$ to take account of the contribution from the late time decay of $N_{2}$ in comparison with the original one in eq. (15). Change of the $Y_{N_{1}}$ evolution caused by imposing the production of the required lepton number asymmetry can be found from the solutions.
@@
Their solutions for the same values of $M_{1}$ and $(\lambda_{3},\lambda_{4})$ used in Fig. 1 are plotted as functions of $z_{2}$ in the left panel of Fig. 2. We use the values of $M_{\eta_{R}}$, $M_{2}$ and $M_{3}$ given in (14) with $\delta_{2}=3\times 10^{-5}$ and $\delta_{3}=10^{-9}$ which are adopted referring to the above discussion. The figure shows that the lepton number asymmetry $|Y_{L}|$ plotted by an orange line is generated to converge to a final value at $z_{2}\sim 10$, which is larger compared with a required range to explain the baryon number asymmetry shown by dotted yellow lines. Since the resulting lepton number asymmetry takes a value $Y_{L}\simeq\varepsilon Y_{N_{2}}^{\rm eq}$ where $Y_{N_{2}}^{\rm eq}$ is the one of the relativistic particles, the washout of the generated lepton number asymmetry is found to be sufficiently suppressed as a result of the late time decay of $N_{2}$ as expected from the above discussion.
@@
Fig. 2  Evolution of the lepton number asymmetry $|Y_{L}|$ and the $N_{1}$ number density $Y_{N_{1}}$ in the comoving volume. $|Y_{L}|$ is plotted by a orange line, which is caused by the out-of-equilibrium decay of $N_{2}$. $Y_{N_{2}}^{\rm eq}$ should be noted to coincide with $Y_{\eta_{R}}^{\rm eq}$ since $M_{2}\simeq M_{\eta_{R}}$ is assumed. Yellow dashed horizontal lines represent a region of $|Y_{L}|$ to explain the required baryon number asymmetry. In the left panel, $Y_{N_{1}}$ is plotted by red lines for the same $M_{1}$ as the ones in the left panel of Fig. 1. Only $h_{1}$ is tuned to realized the present DM abundance which is represented by black solid, dashed and dotted lines for each $M_{1}$. Other parameters are fixed to the same values as the ones in Fig. 1. In the right panel, $(\lambda_{3},\lambda_{4})$ is changed to $(2.5,-1.0)$ from the ones used in the left panel keeping the $h_{1}$ value unchanged. While $Y_{N_{1}}$ becomes a little bit smaller than the required value $Y_{N_{1}}^{\infty}$ in $M_{1}=150$ GeV and 1.5 GeV, $Y_{N_{1}}+Y_{\eta_{R}}=Y_{N_{1}}^{\infty}$ is realized for $M_{1}=1.5$ TeV at $z_{2}=300$. It means that $Y_{N_{1}}=Y_{N_{1}}^{\infty}$ is obtained finally through the $\eta_{R}$ decay.

### v2

@@ 2.2 Relic abundance of right-handed neutrino DM
Fig. 1   Left panel: Evolution of the number density of $N_{1}$ and $\eta_{R}$ in the comoving volume. They are plotted with red lines for three values of $M_{1}$, 1.5 TeV (solid), 150 GeV (dashed) and 1.5 GeV (dotted). $h_{1}$ is fixed at $7.8\times 10^{-9}$, $4.5\times 10^{-8}$ and $7.0\times 10^{-6}$ for each $M_{1}$, respectively. $M_{\eta_{R}}=3$ TeV and $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$ are assumed. Horizontal black lines represent $Y_{N_{1}}^{\infty}$ values required to realize the present DM relic density for each value of $M_{1}$. While $Y_{\eta_{R}}$ for $M_{1}=1.5$ TeV deviates from $Y_{\eta_{R}}^{\rm eq}$, $Y_{\eta_{R}}$ in other two cases coincides with $Y_{\eta_{R}}^{\rm eq}$ since the $\eta_{R}$ decay is already in the equilibrium at a temperature higher than $T_{\eta}$. Right panel: Contours of $Y_{\eta_{R}}(T_{\eta})$ for $M_{\eta_{R}}=3$ TeV in the $(\lambda_{3},\lambda_{4})$ plane. Each red line represents $Y_{\eta_{R}}(T_{\eta})=10^{-10.5},10^{-11},10^{-11.5},10^{-11.75}$, and $10^{-12}$ outward. A dashed line represents $\lambda_{3}+\lambda_{4}+\lambda_{5}=-\sqrt{\lambda_{1}\lambda_{2}}$ in the case of $\lambda_{2}=1.0$, and an upper region of it satisfies the potential stability. No constraint comes from direct detection experiments on $\lambda_{3,4}$ since DM is $N_{1}$.
@@
In order to show this observation concretely, we solve the Boltzmann equations for $Y_{N_{1}}$ and $Y_{\eta_{R}}$ using the parameters given in eq. (14) and $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$. Since $M_{1}$ and $h_{1}$ can be free from neutrino mass determination as addressed below eq. (14), we plot the solutions in the left panel for the cases $(M_{1},h_{1})=(1.5~{\rm TeV},7.8\times 10^{-9})$, $(150~{\rm GeV},4.5\times 10^{-8})$, and $(1.5~{\rm GeV},7.0\times 10^{-6})$ where $h_{1}$ is tuned to give the required DM abundance. In this panel, $Y_{N_{1}}^{\infty}$ required to explain the DM abundance is shown by horizontal black lines for each case.
In the latter two cases, $T_{c}>T_{\eta}$ is satisfied and the required $Y_{N_{1}}^{\infty}$ is found to be realized by the pure freezein due to the $\eta_{R}$ decay in the thermal equilibrium through a relation $Y_{N_{1}}^{\infty}=Y_{N_{1}}(T_{c})=Y_{\eta_{R}}^{\rm eq}(T_{c})$. $Y_{\eta_{R}}$ does not deviate from the equilibrium one there. The freezeout of the $\eta_{R}$ coannihilation in a case of $M_{\eta_{R}}=3$ TeV is shown in the right panel, in which contours for $Y_{\eta_{R}}^{\rm eq}(T_{\eta})=10^{-10.5},10^{-11},10^{-11.5},10^{-11.75}$, and $10^{-12}$ are plotted in the $(\lambda_{3},\lambda_{4})$ plane. Since $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$ is assumed in the left panel, $Y_{\eta_{R}}^{\rm eq}(T_{\eta})$ for it is smaller than the required $Y_{N_{1}}^{\infty}$ from the right panel. The freezeout of the $\eta_{R}$ coannihilation is irrelevant to the estimation of $Y_{N_{1}}^{\infty}$.
As an extreme case for this kind of examples, we may consider $N_{1}$ with keV scale mass. We can confirm that the required relic DM abundance is obtained by tuning $h_{1}$ suitably, for example, $h_{1}=3.7\times 10^{-3}$ for $M_{1}=5$ keV. We note that the values of $h_{2,3}$ in eq. (14) are not affected even if $h_{1}$ takes the same order value as $h_{2,3}$ as long as $M_{1}$ is much smaller than $M_{\eta_{R}}$ as the present case. It is caused by the fact that $\Lambda_{1}$ has a strong suppression factor of $O(M_{1}/M_{\eta_{R}})$ compared with $\Lambda_{2,3}$. Thus, $N_{1}$ with keV scale mass is allowed to have an acceptable relic abundance as DM for $h_{2,3}$ given in eq. (14) without contradicting with the neutrino oscillation data.
@@
In the case of $M_{1}=1.5$ TeV, the freezeout of the $\eta_{R}$ coannihilation is relevant to the estimation of $Y_{N_{1}}$. Although $Y_{N_{1}}(T_{c})$ is smaller than the required value, the late time decay of the frozen $\eta_{R}$ contributes to $Y_{N_{1}}$ and makes it the required value $Y_{N_{1}}^{\infty}$ as found from the figure. Since the stability of the potential $V$ in eq. (1) requires $\lambda_{3,4}$ should satisfy the condition (2), an allowed region of $(\lambda_{3},\lambda_{4})$ is a part over the dashed line in the right panel. If $\lambda_{3}$ and $\lambda_{4}$ are fixed to suitable values in this region, the required $Y_{N_{1}}^{\infty}$ can be obtained. Here, we should note that there is no constraint on the values of $\lambda_{3}$ and $\lambda_{4}$ from direct detection experiments since $N_{1}$ cannot interact with nuclei. It is a crucial difference from the case where $\eta_{R}$ is DM. The coupling constants $\lambda_{3}$ and $\lambda_{4}$ could be also constrained through the correction to the Higgs mass caused by them. This constraint is reported in [31] to be satisfied for $M_{\eta_{R}}\leq 4.4$ TeV.
@@
In the above study, we do not take into account the effect caused by the simultaneous explanation of the baryon number asymmetry in the Universe through leptogenesis which is based on the $N_{2}$ decay. Since it could produce $\eta_{R}$ at late time, the $N_{1}$ abundance could be affected by it if the $N_{2}$ decay occurs near the freezeout of the coannihilation of $\eta_{R}$. We discuss this point in the next part.
@@ 3 Consistency with leptogenesis
In the present model, the baryon number asymmetry is considered to be produced through leptogenesis caused by the out-of-equilibrium decay of $N_{2}$. This requires that $N_{2}$ should be in the thermal bath and its decay to $\eta_{R}$ occurs in a $CP$ asymmetric way at a higher temperature than 100 GeV where sphaleron interaction is in the equilibrium. Values of $h_{2,3}$ given in eq. (14) are large enough to produce $N_{2,3}$ in the thermal bath as relativistic particles through the inverse decay and scatterings caused by them as long as reheating temperature is larger than $M_{3}$. The $CP$ asymmetry in the decay $N_{2}\rightarrow\ell_{\alpha}\eta^{\dagger}$ can be caused through an interference between a tree diagram and a one-loop diagram mediated by $N_{1}$ and $N_{3}$.
This $CP$ asymmetry is estimated through the formula
@@
where $M_{2}\simeq M_{\eta_{R}}$ is taken into account. Since $h_{1}$ is supposed to be very small for the DM explanation as discussed in the previous part, the dominant contribution to this $\varepsilon$ is found to be caused by $N_{3}$ and it can be estimated by using eqs. (7) and (12) asfffAs addressed in the previous part, it should be noted that the existence of both a zero mass eigenvalue and Majorana phases of the right-handed neutrino masses is required to guarantee a non-zero $\varepsilon$ when neutrino Yukawa couplings $h_{\alpha k}$ are real [18].
@@
where parameters in eq. (14) are used and values of $\sin\theta_{ij}$ and $\alpha_{ij}$ are fixed to the ones written below eq. (14). In the second line, $\sin(\gamma_{2}-\gamma_{3})=1$ is assumed. If a smaller $|\lambda_{5}|$ is assumed in eq. (14), Yukawa couplings $h_{2,3}$ take larger values. In that case, $\varepsilon$ takes a larger value but the out-of-equilibrium decay of $N_{2}$ becomes harder.
The $N_{2}$ decay at temperature $T\geq T_{D}$ is in the equilibrium if $T_{D}$ is defined by $\Gamma_{N_{2}}=H(T_{D})$. The $N_{2}$ decay width $\Gamma_{N_{2}}$ can be expressed as
@@
If the washout processes of the generated lepton number asymmetry are frozen out at temperature $T_{F}$, the lepton number asymmetry caused by the out-of-equilibrium decay of $N_{2}$ could escape the washout in the case $T_{D}<T_{F}$. Such a situation could happen if the $N_{2}$ decay is forced to be delayed by making $\Gamma_{N_{2}}$ smaller.
Although $h_{2}$ fixed by the neutrino oscillation data cannot realize such a situation for a TeV scale $M_{2}$, it could be realized through an additional suppression factor for $\Gamma_{N_{2}}$. It is caused by the phase space suppression due to the strict degeneracy between $M_{2}$ and $M_{\eta_{R}}$ [20]. In fact, if their degeneracy is expressed as $M_{\eta_{R}}=(1-\delta_{2})M_{2}$ with $\delta_{2}\ll 1$, $\Gamma_{N_{2}}$ is found to be rewritten with an additional suppression factor $\delta_{2}$ as
$\Gamma_{N_{2}}\simeq\frac{h_{2}^{2}\delta_{2}^{2}}{4\pi}M_{2}=2.9\times 10^{-12}\left(\frac{M_{2}}{3~{\rm TeV}}\right)\left(\frac{\delta_{2}}{10^{-4}}\right)^{2}~{\rm GeV}.$ (24)
If we assume $\delta_{2}<10^{-4}$ at $M_{2}=3$ TeV, $\Gamma_{N_{2}}<H(T)$ is found to be satisfied for $T~{{}^{<}_{\sim}}~M_{2}$ and $N_{2}$ is expected to be out-of-equilibrium there. Thus, if the condition $T_{D}<T_{F}$ is fulfilled, the generated lepton number asymmetry $Y_{L}$ defined as $Y_{L}\equiv Y_{\ell}-Y_{\bar{\ell}}$ could be estimated as $Y_{L}=\varepsilon Y_{N_{2}}^{\rm eq}$ where $Y_{N_{2}}^{\rm eq}$ stands for the equilibrium value of $N_{2}$ as the relativistic particle. In that case, lepton number asymmetry $Y_{L}=O(10^{-10})$ required for the explanation of baryon number asymmetry in the Universe is expected to be obtained if $\varepsilon$ takes a value of $O(10^{-7})$ in eq. (22). Unfortunately, since $\delta_{2}\ll 1$ also suppresses $\varepsilon$ heavily as found from eq.(21), this idea cannot work in the present case.
@@
It is crucial to know what value of $\delta_{2}$ can make the condition $100~{\rm GeV}<T_{D}<T_{F}$ be satisfied for the parameters given in eq. (14). By calculating the relevant reaction rates of washout processes, we find that the freezeout of the $\eta_{R}$ coannihilation occurs for $T_{F}~{{}^{>}_{\sim}}~0.1M_{2}$. Thus, $T_{D}<T_{F}$ is guaranteed if $\delta_{2}<3\times 10^{-5}$ is satisfied. This suggests that the $CP$ asymmetry $\varepsilon$ can take a value of $O(10^{-7})$ in the case $\delta_{3}=O(10^{-9})$, which makes successful leptogenesis possible as addressed below eq. (24). However, if the number density of $N_{2}$ does not decrease to the required $Y_{N_{1}}^{\infty}$ at the temperature where the $\eta_{R}$ coannihilation freezes out, the $\eta_{R}$ produced through the $N_{2}$ decay could contribute to the $N_{1}$ relic density. Thus, the estimation of the $N_{1}$ relic abundance in the previous section cannot naively applied to this scheme for leptogenesis. The effect of $\eta_{R}$ produced in the late time $N_{2}$ decay needs to be taken into account in the estimation of the $N_{1}$ relic abundance.
@@
$\frac{dY_{N_{2}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\Big\{\gamma^{S}(N_{2}\ell)+\gamma^{S}(N_{2}\eta)\Big\}\left(\frac{Y_{N_{2}}^{2}}{Y_{N_{2}}^{{\rm eq}2}}-1\right)+\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)\right],$
$\frac{dY_{L}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[-\varepsilon\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)+\left\{\sum_{k=2,3}\gamma^{D}(N_{k})+2(\gamma^{S}(\ell\eta^{\dagger})+\gamma^{S}(\ell\ell))\right\}\frac{Y_{L}}{Y_{\ell}^{\rm eq}}\right],$
$\frac{dY_{\eta_{R}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\Big\{2\gamma^{S}(\eta_{R}\eta_{R})+\gamma^{D}(\eta_{R})\Big\}\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right)-\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)\right],$
@@
where $\gamma^{S}({N_{2}\ell})$ and $\gamma^{S}(N_{2}\eta)$ are reaction densities of the scattering $N_{3}\ell\leftrightarrow N_{2}\ell$ and $N_{3}\eta\leftrightarrow N_{2}\eta$ for the $N_{2}$ production, and $\gamma^{S}(\ell\eta^{\dagger})$ and $\gamma^{S}(\ell\ell)$ are the ones for the lepton number violating scattering $\ell\eta^{\dagger}\leftrightarrow\bar{\ell}\eta$ and $\ell\ell\leftrightarrow\eta\eta$. A new term is introduced to the equation for $Y_{\eta_{R}}$ to take account of the contribution from the late time decay of $N_{2}$ in comparison with the original one in eq. (15). Change of the $Y_{N_{1}}$ evolution caused by imposing the production of the required lepton number asymmetry can be found from these solutions.
@@
Their solutions for the same values of $M_{1}$ and $(\lambda_{3},\lambda_{4})$ used in Fig. 1 are plotted as functions of $z_{2}$ in the left panel of Fig. 2. We use the values of $M_{\eta_{R}}$, $M_{2}$ and $M_{3}$ given in (14) with $\delta_{2}=3\times 10^{-5}$ and $\delta_{3}=10^{-9}$ which are adopted referring to the above discussion. The figure shows that the lepton number asymmetry $|Y_{L}|$ plotted by an orange line converges to a final value at $z_{2}\sim 10$. It is larger compared with a required range to explain the baryon number asymmetry shown by dotted yellow lines. Since the resulting lepton number asymmetry takes a value $Y_{L}\simeq\varepsilon Y_{N_{2}}^{\rm eq}$ where $Y_{N_{2}}^{\rm eq}$ is the one of the relativistic particles, the washout of the generated lepton number asymmetry is found to be sufficiently suppressed as a result of the late time decay of $N_{2}$. It is expected in the above discussion.
@@
Fig. 2  Evolution of the lepton number asymmetry $|Y_{L}|$ and the $N_{1}$ number density $Y_{N_{1}}$ in the comoving volume. $|Y_{L}|$ is plotted by an orange line, which is caused by the out-of-equilibrium decay of $N_{2}$. $Y_{N_{2}}^{\rm eq}$ coincides with $Y_{\eta_{R}}^{\rm eq}$ since $M_{2}\simeq M_{\eta_{R}}$ is assumed. Yellow dashed horizontal lines represent a region of $|Y_{L}|$ to explain the required baryon number asymmetry. In the left panel, $Y_{N_{1}}$ is plotted by red lines for the same $M_{1}$ as the ones in the left panel of Fig. 1. Only $h_{1}$ is tuned to realized the present DM abundance which is represented by black solid, dashed and dotted lines for each $M_{1}$. Other parameters are fixed to the same values as the ones in Fig. 1. In the right panel, $(\lambda_{3},\lambda_{4})$ is changed to (2.5,1.0)(2.5,-1.0) from the ones used in the left panel keeping the $h_{1}$ value unchanged. While $Y_{N_{1}}$ becomes a little bit smaller than the required value $Y_{N_{1}}^{\infty}$ in the cases $M_{1}=150$ GeV and 1.5 GeV, $Y_{N_{1}}+Y_{\eta_{R}}=Y_{N_{1}}^{\infty}$ is realized for $M_{1}=1.5$ TeV at $z_{2}=300$. It means that $Y_{N_{1}}=Y_{N_{1}}^{\infty}$ is obtained finally through the $\eta_{R}$ decay.